ALC deflection low Replace D21 on IF unit from 7.1 to 9.0V zener.
Readjust all ALC pots.
IF #2 OUtput low Bridge R77 position, replace R78 with R77 470 ohms. See
TKC service Bulletin.
RX Noise (OSC CAR leakage) Shorten L31, L33 leads as much as possible.
Increase NB effectiveness Replace L31, L33 for 33 uh to 27uh, with SHORT
RX noise from strong signals Replace L30 10uh on the IF with self
shielding type (new part L40-1035-21).
TX audio distortion Adjust T15 on the IF approx 1/2 turn back, readjust
all ALC pots up.
10 M PLL unlock change: R77 from 6.8 K to 4.7 K ohm, R94 from 8.2 K to
6.8 K ohm on the PLL unit.
TX audio distortion Adjust 10 W bias for 150mA, 100 W bias for 100mA.
Unlocked all bands Check L2, (Q2 B+) in the PLL unit.
FM'ing with DFC See service bullein of 9/10 for L12, L13 change and add
resistor and shield.